Recent Log Entries

2L03-5 Tuner Stuck on Resonance

Log Number: 3997268Logbook(s): Entry Makers: kernsBody: 

Cavity running at 3Mv/m but tuner isn't working. No faults on tuner so it could be a disconnect downstairs that needs to be checked by SRF.

NDX 2L24&2L25 setup change

Log Number: 3997267Logbook(s): Body: 

I turned the NDX2L24&25 unit HV on at 100V. It keeps the still present leakage current in 2L24 gamma channel limited at about 15 pA currently, but allows the radiation measurements at a non-optimal efficiency (about 50-70% from nominal at 1000V). Makes the dose rates visible and relatively measurable. A rather large neutron dose rates were observed (up to about 10 rem/h in NDX2L24 at 10:30 - 11:30 am). Also decreased the integration time from 5 to 2 sec to stay within the dynamic range in 2L24&2L25.

2L09-8 KCCU Dropped to .82 A

Log Number: 3997266Logbook(s): Entry Makers: kernsBody: 

Cavity struggling at 9.5 Mv/m. High GASK (6.1v and no RATN) at 9.5 Mv/m and found KCCU at .82a and filaments changed from 5.6v to 5.2v somehow ? Cavity needs driveline , RFCM output and 2.5w amp checked or replaced. Leaving at max of 8.0 Mv/m for now.

iocsl26rfint restored

Log Number: 3997262Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2022-05-27 09:45:22 The output from the myRestore execution was as follows: iocsl26rfint Rf12Intlk_Analog@2022-05-27 09:45 55(55) restored Rf12Intlk_Slope@2022-05-27 09:45 121(121) restored Pause: 5 second(s) Execute: caput R2QXIARCinit 1 (success) Old : R2QXIARCinit 0 New : R2QXIARCinit 1 system@2022-05-27 09:45 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl26rfint (success) Procedure return status: Success

2L15-4 Vacuum Change with RF

Log Number: 3997261Logbook(s): Entry Makers: kernsBody: 

Found cavity 4 GSET 0 and noticed big change in vacuum as I tuned it in and got it to ODVH of 9.8 Mv/m. Three was running at 9.0 Mv/m already. Vacuum changed from 5.1413e-10 to settling out to 8.17867e-10 after 4 settled in. They will trip if turned on at the same time and cav. 4 will have to be ramped slowly back up until vacuum corrected. Will leave plots running to see if it gets any better.

MQD1L04 Trim Card Swap

Log Number: 3997260Logbook(s): Entry Makers: bjacobBody: 

MQD1L04 (Trim Rack NL01B02 Ch. 7) current read back was reading 0.014 A above setpoint which was enough to give mismatch alarms. Tried cycling and manually adjusting set point to get the readback to agree but no luck. Went to the rack to reseat the card but the readback was still 0.014 above the setpoint. Swapped out the card and now the readback matches the setpoint.

iocnl1 restored

Log Number: 3997257Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2022-05-27 11:54:47 The output from the myRestore execution was as follows: iocnl1 <NOTE> Rf6Cryo: Provided macro *CAV* is never used <NOTE> Rf6Ops: Provided macro *IOC* is never used Rf6Cryo Execute: burtwb -f /cs/opshome/burt/REBOOT/rf/snap/iocnl1heatdisvINIT.snap (success) Execute: burtwb -f /cs/opshome/burt/REBOOT/rf/snap/iocnl1heatmodemanual.snap (success) Pause: 8 second(s) Rf06Cryo_heatset@2022-05-27 11:54 6(6) restored Pause: 8 second(s) Rf06Cryo_heatmode@2022-05-27 11:54

iocnl1 restored

Log Number: 3997256Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2022-05-27 11:54:47 The output from the myRestore execution was as follows: iocnl1 <NOTE> Rf6Cryo: Provided macro *CAV* is never used <NOTE> Rf6Ops: Provided macro *IOC* is never used Rf6Cryo Execute: burtwb -f /cs/opshome/burt/REBOOT/rf/snap/iocnl1heatdisvINIT.snap (success) Execute: burtwb -f /cs/opshome/burt/REBOOT/rf/snap/iocnl1heatmodemanual.snap (success) Pause: 8 second(s) Rf06Cryo_heatset@2022-05-27 11:54 6(6) restored Pause: 8 second(s) Rf06Cryo_heatmode@2022-05-27 11:54

COMPLETED Replace Negative BPS in E202B12

Log Number: 3997253Logbook(s): Body: 

With Radcon help the bad power module was removed and a spare inserted. Brought rack up and cycled all cards, all appears well. Suspect a fan may have failed in the bad supply.
It looks like all the expected resistance values updated back to some default value (1 Ohm for a BT, 0.5 Ohm for a BC) causing some cards to throw resistance alarms. I remember this happened after coming up from the last down due to an old restore file I believe. I'll ask Gary to take a look at this.

iocnl1 restoration failed

Log Number: 3997252Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2022-05-27 11:54:47 The output from the myRestore execution was as follows: Failed to set restore status indicator iocnl1 <NOTE> Rf6Cryo: Provided macro *CAV* is never used <NOTE> Rf6Ops: Provided macro *IOC* is never used Rf6Cryo Execute: burtwb -f /cs/opshome/burt/REBOOT/rf/snap/iocnl1heatdisvINIT.snap (failure) channel >R14XHTDH.DISV< not connected, no ca_array_put() channel >R13XHTDH.DISV< not connected, no ca_array_put() channel >R12XHTDH.DISV< not connected, no ca_array_

COMMENT ON Power Up and Perform HCO on 10 Amp Trim System from Spring 2022 SAD

Log Number: 3997249Logbook(s): Body: 

The terminal leads @ MDB1L24H and MQB1L24 where swap @
the terminal block. Swap the leads to their correct position and did
polarity and Inductance measurement on both magnets.

MDB1L25V was showing a Open @ the magnet, tightened the
terminal connection and moved the cable around a bit. The magnet
appears to be working now. The magnet cable itself appears to
have been repair at some point and there could be a bad
connection @ the repair portion of the cable... We will continue to
monitor this magnet.

Diggs, Kissinger

Pages